Laplace's equation holds that Uxx+Uyy=0,verify that the second derivative of the following equations are laphace's equation.

1.u=in√x²+y².

2.u=x²-y².


1
Expert's answer
2021-06-29T11:10:40-0400

Part 1

"u=\ln \sqrt{x²+y²}"

"u=\frac{1}{2} \ln (x²+y²) \implies e^{2u}= x^2+y^2"

Differentiating partially with respect to x and y

"2e^{2u}.\frac{∂u}{∂x}=2x \implies U_x= xe^{-2u}"

"\therefore U_{xx}=e^{-2u}-2xe^{-2u}. U_x= e^{-2u}[1-2xU_x]= e^{-2u}[1-2x^2e^{-2u}]"

Similary

"2e^{2u}. U_y=2y \implies U_y=ye^{-2u}"

"\therefore U_{yy}=e^{-2u}-2ye^{-2u}. U_y= e^{-2u}[1-2yU_y]= e^{-2u}[1-2y^2e^{-2u}]"

"U_{xx}+U_{yy}=e^{-2u}[2-2e^{-2u}(x^2+y^2)]"

"U_{xx}+U_{yy}=e^{-2u}[2-2e^{-2u}*e^{2u}]"

"U_{xx}+U_{yy}=e^{-2u}[2-2]"

"U_{xx}+U_{yy}=0"

Thus "u=\ln \sqrt{x²+y²}" is the laphace's equation.


Part 2

"u=x\u00b2-y\u00b2"

"U_x=2x; U_{xx}= 2"

"U_y=-2y; U_{yy}= -2"

"U_{xx}+U_{yy}=2+(-2)"

"U_{xx}+U_{yy}=0"

Thus "u={x\u00b2-y\u00b2}" is the laphace's equation.